OpenStudy (immings):

Not looking for answer, just walk through of which double angle formula to use. SinX=Cos2X

OpenStudy (turingtest):

the one that leaves everything in terms of sine and a constant

OpenStudy (immings):

cos^2x-sin^2x?

OpenStudy (turingtest):

no, that one has a cosine in it...

OpenStudy (immings):

oh soorry dyslexic, thought it said and cosine. so 1-2sin^2x

OpenStudy (turingtest):

yeah :)

OpenStudy (immings):

sinx / 2sin^2x = 1 1 / 2sinx = 1

OpenStudy (turingtest):

whoa there? what did you do?

OpenStudy (turingtest):

sinx=1-sin^2x

OpenStudy (immings):

moved 2sin^2 x to the right side

OpenStudy (immings):

left*

OpenStudy (turingtest):

how, by division? if so you did the algebra wrong

OpenStudy (immings):

oh, I see. I forgot to type the negative sign

OpenStudy (turingtest):

you divided?

OpenStudy (immings):

yes

OpenStudy (turingtest):

if you divide both sides by 2sin^2x you get cscx=csc^2x-1 that doesn't really help us... try addition to get the everything on the left hand side

terenzreignz (terenzreignz):

Might I suggest removing confusion by replacing all instances of sin x with, say, u, to make it look more algebra-y? :)

OpenStudy (turingtest):

sorry, you would get 1/2cscx=1/2csc^2x-1

OpenStudy (immings):

That's not how it is shown from previous examples. there's no addition

OpenStudy (turingtest):

@terenzreignz yeah I was gonna get to that when we got it =0 to show that it looks quadratic, but you could of course do it now as well :)

OpenStudy (turingtest):

@immings Previous example does not matter. all problem,s are different. I'm telling you to us addition (and subtraction) to get everything on the left hand side

OpenStudy (immings):